Can you tell me the steps in solving this:
If the base of a triangle is 11 cm and the area is 242 cm2, what is the height?
(the 2 after the cm is a squared sign)
I know the formulas for area but I don't know how to get the height using the base and area

Answer by scott8148(5880)

the area is one half times base times height

A = 1/2 * b * h

multiplying by 2 ___ 2A = b * h

dividing by b ___ 2A / b = h